OpenStudy (anonymous):

y varies inversely with x, and x = 120 when y = 90. Find the inverse variation equation and use it to find the value of x when y = 30. 210 40 360 22.5

OpenStudy (anonymous):

\[\huge y=\frac{k}{x}\rightarrow 90=\frac{k}{120} \] solve for k then find x with the new y value....

OpenStudy (anonymous):

360

OpenStudy (anonymous):

\[90/30=120^{-1}/x ^{-1}\] x=360

OpenStudy (anonymous):

thanks(: for the help